# Advancing African-Accented English Speech Recognition: Epistemic Uncertainty-Driven Data Selection for Generalizable ASR Models

### v4

@@ 1 Introduction
Active Learning (AL) is a machine learning (ML) technique that leverages epistemic uncertainty (EU) to select the most informative data points from an unlabeled dataset for labeling, thereby improving model performance with fewer training instances. Common types of AL include Deep Bayesian Active Learning (DBAL) [19, 23] and Adversarial Active Learning (AAL) [15]. AAL selects examples likely to be misclassified by the current model, refining it iteratively by challenging it with difficult cases to enhance robustness. The core-set approach (CSA) [42] is also related, as it selects a subset of training data to ensure a model trained on this subset performs comparably to one trained on the full dataset, addressing scalability and efficiency.
Acquisition functions (AFs) in AL select the most informative samples from an unlabeled dataset for labeling. Key AFs include uncertainty sampling (US) [31], Bayesian Active Learning by Disagreement (BALD) [19], and BatchBALD [26]. US targets data points with the highest model uncertainty. BALD maximizes the mutual information between model parameters and predictions. BatchBALD is an extension of BALD which selects multiple samples simultaneously but may choose redundant points. US is the least computationally expensive, ideal for efficient data labeling.
In this work, we leverage and combine DBAL, AAL, US, and CSA in the following way (in order):
  1. 1.
    we integrate the CSA by leveraging smaller training subsets ($\sim 45\%$ smaller than the full available training sets)
  2. 2.
    we use DBAL with MC-Dropout, to apply dropout during training and inference to estimate Bayesian posterior distribution. This allows us to practically and efficiently estimate EU in the models used [19] (see section 3.2 for more details).
  3. 3.
    we use the estimated EU, and integrate the idea of AAL by using the US acquisition function.
@@ 2.2 Active Learning
AL aims at reducing the number of training examples to be labeled by automatically processing the unlabeled examples and then selecting the most informative ones concerning a given cost function for a human to label. It is particularly effective when labeled data is scarce or expensive, optimizing the learning process by focusing on samples that most improve the model performance and generalization [43, 19]. Several works have demonstrated its effectiveness and efficiency. An AL setup involves an unlabeled dataset $\mathcal{D}_{\mathrm{pool}}=\{\mathbf{x}_{i}\}_{i=1}^{n_{\mathrm{pool}}}$, a labeled training set $\mathcal{D}_{\mathrm{train}}=\{\mathbf{x}_{i},\mathrm{y}_{i}\}_{i=1}^{n_{\mathrm{train}}}$, and a predictive model with likelihood $p_{w}(y|x)$ parameterized by $w\sim p(W|\mathcal{D}_{\mathrm{train}})$. The setup assumes the presence of an oracle to provide predictions $\mathrm{y}$ for all $x_{i}\in\mathcal{D}_{\mathrm{pool}}$. After training, a batch of data $\{\mathbf{x}^{*}_{i}\}_{i=1}^{b}$ is selected from $\mathcal{D}_{\mathrm{pool}}$ based on its EU.
@@ 3.1 Datasets
Item Train ($\mathcal{D}_{\mathrm{train}}^{*}$ Train) Dev Test AL Top-k

### v5

@@ 1 Introduction
To further enhance the efficiency and effectiveness of model adaptation, we employ Active Learning (AL) techniques. AL leverages epistemic uncertainty to select the most informative data points from an unlabeled dataset for labeling, thereby improving model performance with fewer training instances. Common types of AL include Deep Bayesian Active Learning (DBAL) [19, 23] and Adversarial Active Learning (AAL) [15]. AAL selects examples likely to be misclassified by the current model, refining it iteratively by challenging it with difficult cases to enhance robustness. The core-set approach (CSA) [42] is also related, as it selects a subset of training data to ensure a model trained on this subset performs comparably to one trained on the full dataset, addressing scalability and efficiency. A critical component of AL is the acquisition function (AF), which determines the most informative samples from an unlabeled dataset for labeling. Key AFs include uncertainty sampling (US) [31], Bayesian Active Learning by Disagreement (BALD) [19], and BatchBALD [26]. US targets data points with the highest model uncertainty. BALD maximizes the mutual information between model parameters and predictions. BatchBALD is an extension of BALD that selects multiple samples simultaneously but may choose redundant points. US is the least computationally expensive, making it ideal for efficient data labeling.
In this work, we leverage and combine DBAL, AAL, US, and CSA in the following way (in order): First, we integrate the CSA by leveraging smaller training subsets ($\sim 45\%$ smaller than the full available training sets). Second, we use DBAL with MC-Dropout, to apply dropout during training and inference to estimate Bayesian posterior distribution. This allows us to practically and efficiently estimate EU in the models used [19] (see section 3.2 for more details). Third, we use the estimated EU and integrate the idea of AAL by using the US acquisition function.
@@ 2.2 Active Learning
AL aims to reduce the number of labeled training examples by automatically processing the unlabeled examples and then selecting the most informative ones concerning a given cost function for a human to label. It is particularly effective when labeled data is scarce or expensive, optimizing the learning process by focusing on samples that most improve the model performance and generalization [43, 19]. Several works have demonstrated its effectiveness and efficiency. An AL setup involves an unlabeled dataset $\mathcal{D}_{\mathrm{pool}}=\{\mathbf{x}_{i}\}_{i=1}^{n_{\mathrm{pool}}}$, a labeled training set $\mathcal{D}_{\mathrm{train}}=\{\mathbf{x}_{i},\mathrm{y}_{i}\}_{i=1}^{n_{\mathrm{train}}}$, and a predictive model with likelihood $p_{w}(y|x)$ parameterized by $w\sim p(W|\mathcal{D}_{\mathrm{train}})$. The setup assumes the presence of an oracle to provide predictions $\mathrm{y}$ for all $x_{i}\in\mathcal{D}_{\mathrm{pool}}$. After training, a batch of data $\{\mathbf{x}^{*}_{i}\}_{i=1}^{b}$ is selected from $\mathcal{D}_{\mathrm{pool}}$ based on its EU.
@@ 3.1 Datasets
Item Train ($\mathcal{D}_{\mathrm{train}}^{*}$) Dev Test AL Top-k